OpenStudy (aaronandyson):

The number of sides of two regular polygons is 5:4 nd the difference between them is 9 degrees .Find the no.of sides of the regular polygon.

OpenStudy (aaronandyson):

@Qwertty123 @ganeshie8 @rebeccaxhawaii

OpenStudy (qwertty123):

Let N be the number of sides of polygon with fewer sides. The other polygon then has 5N/4 sides. 180 - 360/N - [180 - 360*0.8/N] = -9 -360/N + 288/N = -9 72/N = 9 N = 8 The other polygon has 10 sides. The internal angles are 135 and 144 degrees.

OpenStudy (aaronandyson):

pl use an equation editor..